# Techniques for improved statistical convergence in quantification of eddy diffusivity moments

### v1

@@ Abstract
While recent approaches, such as the macroscopic forcing method (MFM) or Green’s function-based approaches, can be used to compute Reynolds-averaged Navier–Stokes closure operators using forced direct numerical simulations, MFM can also be used to directly compute moments of the effective nonlocal and anisotropic eddy diffusivities. The low-order spatial and temporal moments contain limited information about the eddy diffusivity but are often sufficient for quantification and modeling of nonlocal and anisotropic effects. However, when using MFM to compute eddy diffusivity moments, the statistical convergence can be slow for higher-order moments. In this work, we demonstrate that using the same direct numerical simulation (DNS) for all forced MFM simulations improves statistical convergence of the eddy diffusivity moments. We present its implementation in conjunction with a decomposition method that handles the MFM forcing semi-analytically and allows for consistent boundary condition treatment, which we develop for both scalar and momentum transport. We demonstrate that for a two-dimensional Rayleigh–Taylor instability case study, when this technique is applied with $\mathcal{O}(100)$ simulations, we achieve a similar quality of convergence as with $\mathcal{O}(1000)$ simulations using the original method. We then demonstrate the impacts of improved convergence on the quantification of the eddy diffusivity.
@@ I Introduction
In MFM, two sets of equations are solved: the donor and receiver equations. The donor equations provide flow fields needed for the solution of the receiver equations, which are macroscopically forced when probing the eddy diffusivity moments. For example, in the case of determining the closure operator for mean scalar transport in natural convection, the donor equations would be the Navier–Stokes equations two-way-coupled with the scalar transport equation, and the receiver equations would be forced scalar transport equations with different macroscopic forcings. Each receiver equation is forced to obtain a specific spatiotemporal moment of the eddy diffusivity kernel. For ease of implementation, it is natural to run simulations for each eddy diffusivity moment separately where each simulation solves its own donor and receiver equations. However, due to numerical differences between processes in the simulations, the repeated numerical solutions to the donor equations may have differences. Due to the chaotic nature of turbulence, this error can amplify and lead to $\mathcal{O}(1)$ difference in numerical solutions to the donor equations. As we show, this instantaneous difference can propagate as error when determining higher-order moments and lead to slow statistical convergence of the eddy diffusivity moments with respect to the number of realizations.
@@
To avoid this slow convergence, an MFM simulation can instead use one donor to service multiple receiver equations. That is, the individual MFM simulations described in the previous paragraph can be combined into one simulation using one donor but solving for multiple sets of receiver equations. Using one donor then prevents statistical error from random processes and reduces the overall computational cost of MFM, since only one donor needs to be solved.
@@
With this method, we can also utilize the decomposition introduced by Liu et al. [8] to semi-analytically treat the MFM forcing; we call this combined single donor and semi-analytical forcing approach the decomposition MFM. The decomposition method alleviates boundary condition issues when the MFM forcing is incompatible with the boundary conditions of the problem. Liu et al. [8] demonstrated the decomposition method for scalar transport in a steady laminar inhomogeneous flow. In this work, we extend the decomposition method to obtain spatiotemporal moments of the eddy diffusivity for general unsteady and chaotic flows as well as momentum transport. We then apply the decomposition MFM to scalar transport in Rayleigh-Taylor (RT) instability to illustrate the acceleration of statistical convergence due to use of a single donor and its compatibility with the proposed decomposition method.
@@
In this work, we begin with scalar transport before generalizing to momentum transport. In Section II, we define the nonlocal and anisotropic eddy diffusivity and its moments. In Section III, we introduce MFM for directly computing the eddy diffusivity moments. In Section IV, we develop the decomposition method for general unsteady and chaotic flows. In Section V, we generalize to momentum transport and the nonlocal and anisotropic eddy viscosity. In Section VI, as an illustrative example, we demonstrate the improved statistical convergence of the eddy diffusivity moments for 2D RT instability and discuss the modeling impacts.
@@ II Problem formulation for scalar transport
[…]
@@ III Eddy diffusivity moments using the macroscopic forcing method (MFM)
where we have omitted the subscripts on the moments, since here we only consider one spatial dimension. To obtain the zeroth-order spatiotemporal moment of the eddy diffusivity, one specifies $\langle c\rangle=x_{1}$ using inverse MFM and solves the forced scalar transport equation in (10). At each time step, the forcing is used to maintain the specified $\langle c\rangle$ while $c$ is free to evolve. Practically, one can first time advance the governing equation without the forcing and solve for an intermediate scalar field, and then add the forcing in a correction step to ensure the scalar field at the next time step has the requisite $\langle c\rangle$ as discussed in [6, 12]. Postprocessing of $-\langle u_{1}^{\prime}c^{\prime}\rangle$ leads to the zeroth moment:
$-\langle u_{1}^{\prime}c^{\prime}\rangle|_{\langle c\rangle=x_{1}}(x_{1},t)=D^{00}(x_{1},t)$ (12)
@@
$-\langle u_{1}^{\prime}c^{\prime}\rangle|_{\langle c\rangle=x_{1}^{2}/2}(x_{1},t)=x_{1}D^{00}(x_{1},t)+D^{10}(x_{1},t)$ (13)
@@
$-\langle u_{1}^{\prime}c^{\prime}\rangle|_{\langle c\rangle=x_{1}^{3}/6}(x_{1},t)=\frac{x_{1}^{2}}{2}D^{00}(x_{1},t)+x_{1}D^{10}(x_{1},t)+D^{20}(x_{1},t),$ (14)
@@
$-\langle u_{1}^{\prime}c^{\prime}\rangle|_{\langle c\rangle=x_{1}t}(x_{1},t)=tD^{00}(x_{1},t)+D^{01}(x_{1},t)$ (15)
@@
$-\langle u_{i}^{\prime}c^{\prime}\rangle(\mathbf{x},t)|_{\langle c\rangle=x_{\alpha}}=D_{i\alpha}^{00}(\mathbf{x},t).$ (16)
@@
$-\langle u_{i}^{\prime}c^{\prime}\rangle(\mathbf{x},t)|_{\langle c\rangle=x_{\alpha}^{2}/2}=x_{\alpha}D_{i\alpha}^{00}(\mathbf{x},t)+D_{i\alpha\alpha}^{10}(\mathbf{x},t),$ (17)
with no summation over $\alpha$ implied. Postprocessing the scalar flux and then subtracting out the contribution from the zeroth-order moment leads to $D_{j\alpha\alpha}^{10}$. Cross components of the first-order spatial moment may be obtained by specifying $\langle c\rangle=x_{\alpha}x_{\beta}$ where $\alpha,\beta=1,2,\text{or }3$ and $\alpha\neq\beta$.
@@
(b) Decomposition method.
Figure 1: Diagrams outlining MFM and the decomposition method in this work, presented in one dimension for simplicity. Superscripts denote variables ($c^{i}$, $\mathscr{c}^{i}$, etc.) and operators ($\mathcal{L}^{i}$) belonging to the receiver equations solved to obtain $D^{i}$.
@@
Figure 2: Root mean square errors (RMSE) of standard (solid lines) and single donor (dotted lines) MFM as percent of maximum magnitude of each eddy diffusivity moment at each time. Errors are computed at 20 realizations with respect to 200 realizations for each method.
@@ III.1 Discussion of statistical error
A natural implementation of MFM for some codes is to use separate sets of donor and receiver equations for each moment to be computed. This configuration is illustrated in Figure 1(a). This approach is more computationally expensive, since the donor equations are solved multiple times, but is easier to implement, since just one set of receiver equations is one-way coupled to each donor equation. Another approach would be to run an MFM simulation using a single donor for multiple receivers, as illustrated in Figure 1(b) (the additional decomposition features of this proposed method will be detailed in a later section). As we show in Appendix A, while mathematically equivalent to using a single donor, using separate donor simulations can result in larger statistical error. Separate donor simulations, even when set up with identical initial conditions, may produce different instantaneous results due to differences in the order of arithmetic operations (e.g., from parallelization), augmented by the chaotic nature of turbulence simulations. The resulting differences can propagate as error in the computation of higher-order eddy diffusivity moments from lower-order moments through Equations (13)-(17), ultimately leading to slow statistical convergence. When statistical errors between donors match (i.e., a single donor is used for multiple receivers), the amplified error is removed, and statistical convergence is accelerated.
@@
To illustrate, Figure 2 shows the statistical errors associated with the standard and single donor methods applied to the 2D RT problem. In 2D RT instability, the flow is averaged over the homogeneous $x_{1}$ direction and over multiple realizations. Details on the RT case study will be covered later in Section VI, but we present these plots here to demonstrate the differences in error between the two methods. When using Equation (13), computation of $D^{01}$ involves multiplication of $D^{00}$ (and therefore its associated statistical error) by $t$, so we expect error to grow substantially at large $\tau$. This is indeed what we observe in the statistical error plots. Overall, across all higher-order moments, the standard MFM exhibits higher error than the single donor method. We show in Appendix A that when a single donor is used, errors from lower-order moments do not propagate to higher-order moments, so all errors scale similarly with time.
IV Decomposition method
@@
In the previous section, we discussed the need for using the same DNS simulation (donor) for all forced MFM simulations (receivers) to reduce statistical error. In conjunction, we now introduce a decomposition method that can be used simultaneously at no additional cost. The decomposition method treats the MFM forcing semi-analytically and was originally developed to address the issue of periodic boundary conditions in a steady laminar problem [8]. Here we extend the decomposition to general unsteady and chaotic problems and to momentum transport in Section V.
@@
As discussed in Liu et al. [8], the term $-\partial/\partial x_{i}\langle u_{i}^{\prime}c^{\prime}\rangle$ is related to the MFM forcing via
@@
The purely mean scalar terms on the right hand side, $\partial\langle c\rangle/\partial t$ and $D_{M}\partial^{2}\langle c\rangle/\partial x_{i}\partial x_{i}$, can be absorbed into the forcing, $s$, while still maintaining the property $s=\langle s\rangle$. In this alternative formulation for inverse MFM, the forcing $s=\langle s\rangle$ is used to maintain $\langle c^{\prime}\rangle=0$. The right hand side term, $u_{i}\partial\langle c\rangle/\partial x_{i}$, is viewed as a source term to the scalar fluctuation equation. One advantage of using Equation (22) rather than Equation (10) for inverse MFM is that only the derivatives of the mean scalar appear rather than the mean scalar itself. The derivatives of the mean scalar are specified analytically and do not necessarily need a mathematically consistent mean scalar field. This point is further discussed in Liu et al. [13], where we take advantage of the forced fluctuation equation to measure various components of the eddy viscosity tensor in turbulent channel flow independently. The other advantage is that now we can further decompose $c^{\prime}(\mathbf{x},t)$. Similar to the expansion in (6), consider the Taylor series expansion of the general solution for $c^{\prime}(\mathbf{x},t)$ in (19) locally about $\mathbf{x}^{\prime}=\mathbf{x}$ and $t=t^{\prime}$:
@@
By substituting this decomposition for $c^{\prime}(\mathbf{x},t)$ into the forced scalar fluctuation equation in (22), we can derive governing equations for $\mathscr{c}^{mn}(\mathbf{x},t)$. As with inverse MFM, we can activate various $\mathscr{c}^{mn}(\mathbf{x},t)$ fields by specifying the mean scalar gradient. The main difference from inverse MFM as detailed in Section III is that the mean scalar gradient and MFM forcing are now treated semi-analytically. For example, specifying the mean scalar gradient as $\partial\langle c\rangle/\partial x_{j}=1$ leads to an equation for $\mathscr{c}^{00}(\mathbf{x},t)$. Specifying the mean scalar gradient as higher-order polynomials leads to higher-order $\mathscr{c}^{mn}(\mathbf{x},t)$, and equations for the lower-order $\mathscr{c}^{mn}(\mathbf{x},t)$ can be analytically subtracted to derive an equation for the desired order of $\mathscr{c}^{mn}(\mathbf{x},t)$. The $\mathscr{c}^{mn}(\mathbf{x},t)$ fields can then be used to compute the moments by multiplying (23) by $-u_{i}^{\prime}$ and averaging:
@@
where the inverse MFM forcing maintains the specified mean scalar gradient by equivalently maintaining $\langle c^{\prime}\rangle(x_{1},t)=\langle\mathscr{c}^{00}\rangle(x_{1},t)=0$. Postprocessing $-\langle u_{1}^{\prime}\mathscr{c}^{00}\rangle(x_{1},t)$ leads to the zeroth-order moment of the eddy diffusivity, $D^{00}(x_{1},t)$.
@@
where we have relabeled the inverse MFM forcing as $s^{01}=s-s^{00}t$. The forcing maintains $\langle\mathscr{c}^{01}\rangle(x_{1},t)=0$. Postprocessing $-\langle u_{1}^{\prime}\mathscr{c}^{01}\rangle(x_{1},t)$ leads to the first-order temporal moment of the eddy diffusivity, $D^{01}(x_{1},t)$. Note the equation for $\mathscr{c}^{01}(\mathbf{x},t)$ is coupled with the equation for $\mathscr{c}^{00}(\mathbf{x},t)$. Generally, higher-order $\mathscr{c}^{mn}(\mathbf{x},t)$ are one-way coupled with lower-order $\mathscr{c}^{mn}(\mathbf{x},t)$.
The cost of using the decomposition method is identical to the cost of MFM. By treating the forcing semi-analytically, the decomposition method can be used for periodic problems where the mean scalar gradient needed for MFM, e.g., $\partial\langle c\rangle/\partial x_{1}=x_{1}$, may be incompatible with periodic boundary conditions. The equations for the decomposed variables satisfy the periodic boundary conditions and all explicit dependence on the coordinate is analytically removed. Moreover, the decomposition method also allows one to probe different directions of the eddy diffusivity independently. For example, consider a 2D problem, where $-\langle u_{i}^{\prime}c^{\prime}\rangle(x_{1},x_{2})$ and $\langle c\rangle(x_{1},x_{2})$. The decomposition method allows one to specify various directions of the mean scalar gradient, e.g., $\partial\langle c\rangle/\partial x_{1}=x_{2}$ and $\partial\langle c\rangle/\partial x_{2}=0$, even when a $\langle c\rangle(x_{1},x_{2})$ that satisfies both desired gradients may not exist. This allows one to probe different directions of the eddy diffusivity, e.g., quantify $D_{11}$ independently from $D_{12}$. However, the resulting closure for $-\langle u_{i}^{\prime}c\rangle$ is still mathematically consistent as a linear superposition of the various components of the eddy diffusivity.

### v2

@@ Abstract
While recent approaches, such as the macroscopic forcing method (MFM) or Green’s function-based approaches, can be used to compute Reynolds-averaged Navier–Stokes closure operators using forced direct numerical simulations, MFM can also be used to directly compute moments of the effective nonlocal and anisotropic eddy diffusivities. The low-order spatial and temporal moments contain limited information about the eddy diffusivity but are often sufficient for quantification and modeling of nonlocal and anisotropic effects. However, when using MFM to compute eddy diffusivity moments, the statistical convergence can be slow for higher-order moments. In this work, we demonstrate that using the same direct numerical simulation (DNS) for all forced MFM simulations improves statistical convergence of the eddy diffusivity moments. We present its implementation in conjunction with a decomposition method that handles the MFM forcing semi-analytically and allows for consistent boundary condition treatment, which we develop for both scalar and momentum transport. We demonstrate that for a two-dimensional Rayleigh–Taylor instability case study, using the same DNS for all forced MFM simulations results in convergence with $\mathcal{O}(100)$ simulations rather than $\mathcal{O}(1000)$ simulations. We then demonstrate the impacts of improved convergence on the quantification of the eddy diffusivity.
@@ I Introduction
In MFM, two sets of equations are solved: the donor and receiver equations. The donor equations provide flow fields needed for the solution of the receiver equations, which are macroscopically forced for probing the eddy diffusivity moments. For example, in the case of determining the closure operator for mean scalar transport in natural convection, the donor equations would be the Navier–Stokes equations two-way-coupled with the scalar transport equation, and the receiver equations would be forced scalar transport equations with various macroscopic forcings. Each forcing is used to obtain a specific spatiotemporal moment of the eddy diffusivity kernel. For ease of implementation, it is natural to run simulations for each eddy diffusivity moment separately where each simulation solves its own donor and receiver equations. However, due to numerical differences between processes in the simulations, the repeated numerical solutions of the donor equations may have differences. The chaotic nature of turbulence can amplify these errors and lead to $\mathcal{O}(1)$ differences in the numerical solutions of the donor equations. In this work, we examine how small differences in the donor solutions can affect the statistical error in the eddy diffusivity moments. Moreover, for chaotic problems with limited homogeneous dimensions, statistical convergence of the MFM-measured eddy diffusivity moments must be achieved through averaging many realizations of the flow. We examine the importance of using identical sets of realizations, rather than randomly generating realizations, for computing eddy diffusivity moments. As we show, differences in the donor (whether by design choice or through computational differences) can propagate errors when determining higher-order moments and lead to slow statistical convergence of the eddy diffusivity moments.
@@
To avoid this slow convergence, an MFM simulation can instead use one donor to service multiple receiver equations. That is, the individual MFM simulations described in the previous paragraph can be combined into one simulation using one donor but solving for multiple sets of receiver equations corresponding to multiple spatiotemporal moments of the eddy diffusivity. Using a single donor then prevents statistical error from random processes and reduces the overall computational cost of MFM, since only one donor needs to be solved.
@@
With a single donor approach, we can then also use the decomposition introduced by Liu et al. [8] to semi-analytically treat the MFM forcing. The decomposition method (known as decomposition MFM in this work) alleviates boundary condition issues when the MFM forcing is incompatible with the boundary conditions of the problem. Liu et al. [8] demonstrated the decomposition method for scalar transport in a steady laminar inhomogeneous flow. In this work, we extend the decomposition method to obtain spatiotemporal moments of the eddy diffusivity for general unsteady and chaotic flows as well as momentum transport. We then apply decomposition MFM to scalar transport in Rayleigh-Taylor (RT) instability to illustrate the acceleration of statistical convergence due to use of a single donor and its compatibility with the proposed decomposition method.
@@
In this work, we begin with scalar transport before generalizing to momentum transport. In Section II, we define the nonlocal and anisotropic eddy diffusivity and its moments. In Section III, we introduce MFM for directly computing the eddy diffusivity moments. In Section IV, we develop decomposition MFM for general unsteady and chaotic flows. In Section V, we generalize to momentum transport and the nonlocal and anisotropic eddy viscosity. In Section VI, as an illustrative example, we demonstrate the improved statistical convergence of the eddy diffusivity moments for 2D RT instability and discuss the modeling impacts.
@@ II Problem formulation for scalar transport
where $c(\mathbf{x},t)$ is a passive scalar field, $u_{i}(\mathbf{x},t)$ is an incompressible velocity field, and $D_{M}$ is the molecular diffusivity. The average, which we denote using $\langle\cdot\rangle$, may be defined as an ensemble average if the flow is ergodic, a temporal average if the flow is statistically stationary, and/or a spatial average if the flow is statistically homogeneous. Substitution of the Reynolds decomposition, $c=\langle c\rangle+c^{\prime}$, into the scalar transport equation in (1) and averaging gives
@@ III Eddy diffusivity moments using the macroscopic forcing method (MFM)
where we have omitted the tensor subscripts on the moments, since here we only consider one spatial dimension. To obtain the zeroth-order spatiotemporal moment of the eddy diffusivity, one specifies $\langle c\rangle=x_{1}$ using inverse MFM and solves the forced scalar transport equation in (10). At each time step, the forcing is used to maintain the specified $\langle c\rangle$ while $c$ is free to evolve. Practically, one can first time advance the governing equation without the forcing and solve for an intermediate scalar field, and then add the forcing in a correction step to ensure the scalar field at the next time step has the requisite $\langle c\rangle$ as discussed in [6, 12]. Postprocessing of $\langle u_{1}^{\prime}c^{\prime}\rangle$ leads to the zeroth moment:
$\langle u_{1}^{\prime}c^{\prime}\rangle|_{\langle c\rangle=x_{1}}(x_{1},t)=D^{00}(x_{1},t)$ (12)
@@
$\langle u_{1}^{\prime}c^{\prime}\rangle|_{\langle c\rangle=x_{1}^{2}/2}(x_{1},t)=x_{1}D^{00}(x_{1},t)+D^{10}(x_{1},t)$ (13)
@@
$\langle u_{1}^{\prime}c^{\prime}\rangle|_{\langle c\rangle=x_{1}^{3}/6}(x_{1},t)=\frac{x_{1}^{2}}{2}D^{00}(x_{1},t)+x_{1}D^{10}(x_{1},t)+D^{20}(x_{1},t),$ (14)
@@
$\langle u_{1}^{\prime}c^{\prime}\rangle|_{\langle c\rangle=x_{1}t}(x_{1},t)=tD^{00}(x_{1},t)+D^{01}(x_{1},t)$ (15)
@@
$\langle u_{i}^{\prime}c^{\prime}\rangle(\mathbf{x},t)|_{\langle c\rangle=x_{\alpha}}=D_{i\alpha}^{00}(\mathbf{x},t).$ (16)
@@
$\langle u_{i}^{\prime}c^{\prime}\rangle(\mathbf{x},t)|_{\langle c\rangle=x_{\alpha}^{2}/2}=x_{\alpha}D_{i\alpha}^{00}(\mathbf{x},t)+D_{i\alpha\alpha}^{10}(\mathbf{x},t),$ (17)
with no summation over $\alpha$ implied. Postprocessing the scalar flux and then subtracting out the contribution from the zeroth-order moment leads to $D_{i\alpha\alpha}^{10}$. Cross components of the first-order spatial moment may be obtained by specifying $\langle c\rangle=x_{\alpha}x_{\beta}$ where $\alpha,\beta=1,2,\text{or }3$ and $\alpha\neq\beta$.
@@
(b) MFM using a single donor and decomposition method.
Figure 1: Diagrams outlining MFM implementation with separate donors vs. single donor and the decomposition method in this work, presented in one dimension for simplicity. Superscripts denote variables ($c^{i}$, $\mathscr{c}^{i}$, etc.) and operators ($\mathcal{L}^{i}$) belonging to the receiver equations solved to obtain $D^{i}$.
@@
Figure 2: Root mean square errors (RMSE) of MFM with separate donors (solid lines) vs. single donor (dotted lines) as percent of maximum magnitude of each eddy diffusivity moment at each time. Errors are computed at 20 realizations with respect to 200 realizations for each method.
@@ III.1 Discussion of statistical error
A natural implementation of MFM for some codes is to use separate sets of donor equations (e.g., Navier–Stokes equations coupled with scalar transport equations) and receiver equations (e.g., forced scalar transport equations (Equation 10)) for each moment. This configuration is illustrated in Figure 1(a). This approach is more computationally expensive, since the donor equations are solved multiple times, but is easier to implement, since just one set of receiver equations is one-way coupled to each donor equation.
@@
Another approach would be to run an MFM simulation using a single donor for multiple receivers, as illustrated in Figure 1(b) (the additional decomposition features of this proposed method will be detailed in a later section). As we show in Appendix A, while mathematically equivalent to using a single donor, using separate donor simulations can result in larger statistical error. Separate donor simulations, even when set up with identical initial conditions, may produce different instantaneous results due to differences in the order of arithmetic operations (e.g., from parallelization), augmented by the chaotic nature of turbulence simulations. As another example, for flows that require averaging over many realizations, differences may also arise in the donor simulations due to naive random generation of ensemble members. The resulting differences can propagate as error in the computation of higher-order eddy diffusivity moments from lower-order moments through Equations (13)-(17), ultimately leading to slow statistical convergence. When statistical errors between the donors match, the amplified error is removed, and statistical convergence is accelerated. This can be ensured by using a single donor for multiple receivers. We note that in cases where the numerical solution between separate donors is identical, this approach is not necessary and using separate donors vs. a single donor will give the same result without amplification of error. However, using a single donor has the additional cost-saving advantage of not solving multiple instances of the donor equations and the added benefit of being able to use the decomposition method (detailed in Section IV) for problems where the MFM forcing may be incompatible with the boundary conditions of the problem.
@@
To illustrate a case where separate donors can result in different numerical solutions due to differences in parallel processes amplified by the chaotic nature of turbulence, Figure 2 shows the statistical errors associated with using separate donors vs. a single donor applied to the 2D RT problem in the Ares code (detailed in Section VI.4). In 2D RT instability, the flow is averaged over the homogeneous $x_{1}$ direction and over multiple realizations. Details on the RT case study will be covered later in Section VI, but we present these plots here to demonstrate the differences in error between the two methods. When using Equation (15), computation of $D^{01}$ involves multiplication of $D^{00}$ (and therefore its associated statistical error) by $t$, so we expect error to grow substantially at large $\tau$. This is indeed what we observe in the statistical error plots. Overall, across all higher-order moments, using separate donors exhibits higher error than when using a single donor. We show in Appendix A that when a single donor is used, errors from lower-order moments do not propagate to higher-order moments, so all errors scale similarly with time.
IV Decomposition MFM
@@
In the previous section, we discussed the need for using the same DNS simulation (donor) for all forced MFM simulations (receivers) to reduce statistical error. In conjunction, we now introduce a decomposition method that can be used concurrently at no additional cost when the donor equations and all receiver equations are solved simultaneously. Mathematically equivalent to MFM, the decomposition method treats the MFM forcing semi-analytically and was originally developed to address the issue of periodic boundary conditions in a steady laminar problem [8]. Here we extend the decomposition to general unsteady and chaotic problems and to momentum transport in Section V.
@@
As discussed in Liu et al. [8], the term $\partial/\partial x_{i}\langle u_{i}^{\prime}c^{\prime}\rangle$ is related to the MFM forcing via
@@
The purely mean scalar terms on the right hand side, $\partial\langle c\rangle/\partial t$ and $D_{M}\partial^{2}\langle c\rangle/\partial x_{i}\partial x_{i}$, can be absorbed into the forcing, $s$, while still maintaining the property $s=\langle s\rangle$. In this alternative formulation for inverse MFM, the forcing $s=\langle s\rangle$ is used to maintain $\langle c^{\prime}\rangle=0$. The right hand side term, $u_{i}\partial\langle c\rangle/\partial x_{i}$, is viewed as a source term to the scalar fluctuation equation. One advantage of using Equation (22) rather than Equation (10) for inverse MFM is that only the derivatives of the mean scalar, e.g., $\partial\langle c\rangle/\partial t$ and $\partial\langle c\rangle/\partial x_{i}$, appear rather than the mean scalar, $\langle c\rangle$. The derivatives of the mean scalar are specified analytically and do not necessarily need a mathematically consistent mean scalar field. This point is further discussed in Liu et al. [13], where we take advantage of the forced fluctuation equation to measure various components of the eddy viscosity tensor in turbulent channel flow independently. The other advantage is that now we can further decompose $c^{\prime}(\mathbf{x},t)$. Similar to the expansion in (6), consider the Taylor series expansion of the general solution for $c^{\prime}(\mathbf{x},t)$ in (19) locally about $\mathbf{x}^{\prime}=\mathbf{x}$ and $t=t^{\prime}$:
@@
This expansion is similar to that used in moment-gradient methods for dispersion modeling, such as in the work of Nadim et al. [14]. In that work, the microscale probability density is expressed as an expansion of gradients of the macroscale probability density, analogous to our expansion of the microscopic $c^{\prime}$ in terms of gradients of the macroscopic $\langle c\rangle$. The moment-gradient expansion described by Nadim et al. [14] is used to derive dispersion models; in contrast, MFM does not directly output models but instead utilizes the expansion for numerical measurement of moments.
By substituting the decomposition in (23) for $c^{\prime}(\mathbf{x},t)$ into the forced scalar fluctuation equation in (22), we can derive governing equations for $\mathscr{c}^{mn}(\mathbf{x},t)$. As with inverse MFM, we can activate various $\mathscr{c}^{mn}(\mathbf{x},t)$ fields by specifying the mean scalar gradient. The main difference from inverse MFM as detailed in Section III is that the mean scalar gradient and MFM forcing are now treated semi-analytically. For example, specifying the mean scalar gradient as $\partial\langle c\rangle/\partial x_{j}=1$ leads to an equation for $\mathscr{c}^{00}(\mathbf{x},t)$. Specifying the mean scalar gradient as higher-order polynomials leads to higher-order $\mathscr{c}^{mn}(\mathbf{x},t)$, and equations for the lower-order $\mathscr{c}^{mn}(\mathbf{x},t)$ can be analytically subtracted to derive an equation for the desired order of $\mathscr{c}^{mn}(\mathbf{x},t)$. The $\mathscr{c}^{mn}(\mathbf{x},t)$ fields can then be used to compute the moments by multiplying (23) by $u_{i}^{\prime}$ and averaging:
@@
where the inverse MFM forcing maintains the specified mean scalar gradient by equivalently maintaining $\langle c^{\prime}\rangle(x_{1},t)=\langle\mathscr{c}^{00}\rangle(x_{1},t)=0$. Postprocessing $\langle u_{1}^{\prime}\mathscr{c}^{00}\rangle(x_{1},t)$ leads to the zeroth-order moment of the eddy diffusivity, $D^{00}(x_{1},t)$.
@@
where we have relabeled the inverse MFM forcing as $s^{01}=s-s^{00}t$. The forcing maintains $\langle\mathscr{c}^{01}\rangle(x_{1},t)=0$. Postprocessing $\langle u_{1}^{\prime}\mathscr{c}^{01}\rangle(x_{1},t)$ leads to the first-order temporal moment of the eddy diffusivity, $D^{01}(x_{1},t)$. Note the equation for $\mathscr{c}^{01}(\mathbf{x},t)$ is coupled with the equation for $\mathscr{c}^{00}(\mathbf{x},t)$. Generally, higher-order $\mathscr{c}^{mn}(\mathbf{x},t)$ are one-way coupled with lower-order $\mathscr{c}^{mn}(\mathbf{x},t)$. Since higher-order $\mathscr{c}^{mn}(\mathbf{x},t)$ depend on lower-order $\mathscr{c}^{mn}(\mathbf{x},t)$, all decomposition receiver equations require access to the same donor solution, so the decomposition method naturally uses a single donor simulation.
The cost of using the decomposition method is identical to the cost of MFM. By treating the forcing semi-analytically, the decomposition method can be used for periodic problems where the mean scalar gradient needed for MFM, e.g., $\partial\langle c\rangle/\partial x_{1}=x_{1}$, may be incompatible with periodic boundary conditions. The equations for the decomposed variables satisfy the periodic boundary conditions and all explicit dependence on the coordinate is analytically removed. Moreover, the decomposition method also allows one to probe different directions of the eddy diffusivity independently. For example, consider a 2D problem, where $\langle u_{i}^{\prime}c^{\prime}\rangle(x_{1},x_{2})$ and $\langle c\rangle(x_{1},x_{2})$. The decomposition method allows one to specify various directions of the mean scalar gradient, e.g., $\partial\langle c\rangle/\partial x_{1}=x_{2}$ and $\partial\langle c\rangle/\partial x_{2}=0$, even when a $\langle c\rangle(x_{1},x_{2})$ that satisfies both desired gradients may not exist. This allows one to probe different directions of the eddy diffusivity, e.g., quantify $D_{11}$ independently from $D_{12}$. However, the resulting closure for $\langle u_{i}^{\prime}c\rangle$ is still mathematically consistent as a linear superposition of the various components of the eddy diffusivity.